References

[1] L. Aceto, W.J. Fokkink, R.J. van Glabbeek & A. Ingólfsdóttir (1996): Axiomatizing Prefix Iteration with Silent Steps. Information and Computation 127(1), pp. 26–40, doi:10.1006/inco.

1996.0047. Available at http://theory.stanford.edu/~rvg/abstracts.html#37.

[2] L. Aceto, W.J. Fokkink, R.J. van Glabbeek & A. Ingólfsdóttir (2004): Nested Semantics over Finite Trees are Equationally Hard. Information and Computation 191(2), pp. 203–232, doi:10.

1016/j.ic.2004.02.001. Available at http://theory.stanford.edu/~rvg/abstracts.
html#54.

[3] A. Aliseda, R.J. van Glabbeek & D. Westerståhl, editors (1998): Computing Natural Language. CSLI Lecture Notes 81, Center for the Study of Language and Information, Stanford University.

Available at http://theory.stanford.edu/~rvg/abstracts.html#39.

[4] J.C.M. Baeten & R.J. van Glabbeek (1987): Another look at abstraction in process algebra. Report CS-R8701, CWI, Amsterdam. Available at http://www.cse.unsw.edu.au/~rvg/pub/ another.pdf. Extended abstract in Th. Ottmann, editor: Proceedings 14^{th} International Colloquium on Automata, Languages and Programming, ICALP ’87, Karlsruhe, Germany, July 1987, LNCS 267, Springer, 1987, pp. 84–94, doi: 10.1007/3-540-18088-5_8.

[5] J.C.M. Baeten & R.J. van Glabbeek (1987): Another look at abstraction in process algebra (extended abstract). In Th. Ottmann, editor: Proceedings 14^{th} International Colloquium on Automata, Languages and Programming, ICALP ’87, Karlsruhe, Germany, July 1987, LNCS 267, Springer,

pp. 84–94, doi:10.1007/3-540-18088-5_8. Available at http://www.cse.unsw.edu.au/
~rvg/pub/another-ea.pdf.

[6] J.C.M. Baeten & R.J. van Glabbeek (1987): Merge and termination in process algebra. In K.V. Nori, editor: Proceedings 7^{th}Conference on Foundations of Software Technology and Theoretical Computer Science, Pune, India, December 1987, LNCS 287, Springer, pp. 153–172, doi:10.

1007/3-540-18625-5_49. Available at http://theory.stanford.edu/~rvg/abstracts.
html#4.

[7] J.C.M. Baeten & R.J. van Glabbeek (1989): Abstraction and empty process in process algebra. Fundamenta Informaticae XII, pp. 221–242. Available at http://www.cse.unsw.edu.

au/~rvg/pub/etaepsilon.pdf.

[8] R. Barry, R.J. van Glabbeek & P. Höfner (2020): Formalising the Optimised Link State Routing Protocol. In A. Fehnker & H. Garavel, editors: Proceedings of the 4th Workshop on Models for Formal Analysis of Real Systems, Dublin, Ireland, April 26, 2020, Electronic Proceedings in Theoretical Computer Science 316, Open Publishing Association, pp. 40–71, doi:10.4204/

EPTCS.316.3.

[9] N. Bertrand, L. de Alfaro, R.J. van Glabbeek, C. Palamidessi & N. Yoshida (2021): CONCUR Test-Of-Time Award 2021. In S. Haddad & D. Varacca, editors: Proceedings 32nd International Conference on Concurrency Theory, CONCUR’21, Leibniz International Proceedings in Informatics (LIPIcs) 203, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, doi:10.4230/LIPIcs.

CONCUR.2021.1.

[10] B. Bloom, W.J. Fokkink & R.J. van Glabbeek (2000): Precongruence Formats for Decorated Trace Preorders. In: Proceedings 15^{th}Annual IEEE Symposium on Logic in Computer Science, LICS’00, Santa Barbara, USA, June 2000, IEEE Computer Society Press, pp. 107–118, doi:10.

1109/LICS.2000.855760. Available at http://theory.stanford.edu/~rvg/abstracts.
html#44.

[11] B. Bloom, W.J. Fokkink & R.J. van Glabbeek (2004): Precongruence Formats for Decorated Trace Semantics. Transactions on Computational Logic 5(1), pp. 26–78, doi:10.1145/963927.

963929. Available at http://theory.stanford.edu/~rvg/abstracts.html#48.

[12] T. Bourke, R.J. van Glabbeek & P. Höfner (2014): A mechanized proof of loop freedom of the (untimed) AODV routing protocol. In F. Cassez & J.-F. Raskin, editors: Proceedings 12th International Symposium on Automated Technology for Verification and Analysis, ATVA’14, Sydney, NSW Australia, November 2014, LNCS 8837, Springer, pp. 47–63, doi:10.1007/

978-3-319-11936-6_5. Available at http://arxiv.org/abs/1505.05646.

[13] T. Bourke, R.J. van Glabbeek & P. Höfner (2014): Showing Invariance Compositionally for a Process Algebra for Network Protocols. In G. Klein & R. Gamboa, editors: Proceedings 5th International Conference on Interactive Theorem Proving, ITP’14; held as part of the Vienna Summer of Logic, VSL’14, Vienna, Austria, July 2014, LNCS 8558, Springer, pp. 144–159, doi:10.1007/

978-3-319-08970-6_10. Available at http://arxiv.org/abs/1407.3519.

[14] T. Bourke, R.J. van Glabbeek & P. Höfner (2016): Mechanizing a Process Algebra for Network Protocols. Journal of Automated Reasoning 56(3), pp. 309–341, doi:10.1007/

s10817-015-9358-9. Available at http://arxiv.org/abs/1512.07304.

[15] E. Bres, R.J. van Glabbeek & P. Höfner (2016): A Timed Process Algebra for Wireless Networks with an Application in Routing. Technical Report 9145, NICTA. Available at http://arxiv. org/abs/1606.03663. Extended abstract in P. Thiemann, editor: Programming Languages and Systems: Proceedings 25th European Symposium on Programming, ESOP’16; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’16, Eindhoven, The Netherlands, April 2016, LNCS 9632, Springer, 2016, pp. 95-122.

[16] E. Bres, R.J. van Glabbeek & P. Höfner (2016): A Timed Process Algebra for Wireless Networks with an Application in Routing (extended abstract). In P. Thiemann, editor: Programming Languages and Systems: Proceedings 25th European Symposium on Programming, ESOP’16; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’16, Eindhoven, The Netherlands, April 2016, LNCS 9632, Springer, pp. 95–122, doi:10.1007/

978-3-662-49498-1_5.

[17] N. Busi, R.J. van Glabbeek & R. Gorrieri (1994): Axiomatising ST-bisimulation equivalence. In E.-R. Olderog, editor: Proceedings IFIP TC2 Working Conference on Programming Concepts, Methods and Calculi, San Miniato, Italy, June 1994, IFIP Transactions A-56, North-Holland, pp.

169–188. Available at http://theory.stanford.edu/~rvg/abstracts.html#29.

[18] T. Chen, W.J. Fokkink & R.J. van Glabbeek (2008): On Finite Bases for Weak Semantics: Failures Versus Impossible Futures. Available at http://arxiv.org/abs/0810.4904. Extended abstract in M. Nielsen, A. Kucera, P.B. Miltersen, C. Palamidessi, P. Tuma & F.D. Valencia, editors: Proceedings 35th Conference on Current Trends in Theory and Practice of Computer Science (SOFSEM), Špindlerův Mlýn, Czech Republic, January 24-30, 2009, LNCS 5404, Springer, pp. 167-180.

[19] T. Chen, W.J. Fokkink & R.J. van Glabbeek (2008): Ready to preorder: The case of weak process semantics. Information Processing Letters 109(2), pp. 104–111, doi:10.1016/j.ipl.2008.09.

003. Available at http://theory.stanford.edu/~rvg/abstracts.html#80.

[20] T. Chen, W.J. Fokkink & R.J. van Glabbeek (2009): On Finite Bases for Weak Semantics: Failures Versus Impossible Futures. In M. Nielsen, A. Kucera, P.B. Miltersen, C. Palamidessi, P. Tuma & F.D. Valencia, editors: Proceedings 35th Conference on Current Trends in Theory and Practice of Computer Science (SOFSEM), Špindlerův Mlýn, Czech Republic, January 24-30, 2009, LNCS 5404, Springer, pp. 167–180, doi:10.1007/978-3-540-95891-8_18. Available at http://

theory.stanford.edu/~rvg/abstracts.html#81.

[21] T. Chen, W.J. Fokkink & R.J. van Glabbeek (2015): On the Axiomatizability of Impossible Futures. Logical Methods in Computer Science 11(3):17, doi:10.2168/LMCS-11(3:17)2015.

Available at http://arxiv.org/abs/1505.04985.

[22] I. Czaja, R.J. van Glabbeek & U. Goltz (1992): Interleaving semantics and action refinement with atomic choice. In G. Rozenberg, editor: Advances in Petri Nets 1992, LNCS 609, Springer, pp.

89–107, doi:10.1007/3-540-55610-9_169. Available at http://theory.stanford.edu/
~rvg/abstracts.html#24.

[23] Y. Deng & R.J. van Glabbeek (2010): Characterising Probabilistic Processes Logically. Available at http://arxiv.org/abs/1007.5188. Extended abstract in C.G. Fermüller & A. Voronkov, editors: Proceedings 17th International Conference on Logic for Programming, Artificial Intelligence and Reasoning (LPAR), Yogyakarta, Indonesia, October 2010, LNCS 6397, Springer, 2010, pp. 278-293.

[24] Y. Deng & R.J. van Glabbeek (2010): Characterising Probabilistic Processes Logically (extended abstract). In C.G. Fermüller & A. Voronkov, editors: Proceedings 17th International Conference on Logic for Programming, Artificial Intelligence, and Reasoning (LPAR-17), Yogyakarta, Indonesia, October 10-15, 2010, LNCS 6397, Springer, pp. 278–293, doi:10.1007/

978-3-642-16242-8_20. Available at http://theory.stanford.edu/~rvg/abstracts.
html#86.

[25] Y. Deng, R.J. van Glabbeek, M. Hennessy & C.C. Morgan (2008): Characterising Testing Preorders for Finite Probabilistic Processes. Logical Methods in Computer Science 4(4):4, doi:10.

2168/LMCS-4(4:4)2008. Available at http://theory.stanford.edu/~rvg/abstracts.
html#79.

[26] Y. Deng, R.J. van Glabbeek, M. Hennessy & C.C. Morgan (2009): Testing Finitary Probabilistic Processes (extended abstract). In M. Bravetti & G. Zavattaro, editors: Proceedings 20th International Conference on Concurrency Theory (CONCUR’09), Bologna, Italy, September 1-4, 2009, LNCS 5710, Springer, pp. 274–288, doi:10.1007/978-3-642-04081-8_19. Available at

http://theory.stanford.edu/~rvg/abstracts.html#83.

[27] Y. Deng, R.J. van Glabbeek, M. Hennessy & C.C. Morgan (2011): Real-Reward Testing for Probabilistic Processes (extended abstract). In M. Massink & G. Norman, editors: Proceedings Ninth Workshop on Quantitative Aspects of Programming Languages, Saarbrücken, Germany, April 1-3, 2011, Electronic Proceedings in Theoretical Computer Science 57, Open Publishing Association,

pp. 61–73, doi:10.4204/EPTCS.57.5.

[28] Y. Deng, R.J. van Glabbeek, M. Hennessy & C.C. Morgan (2014): Real-Reward Testing for Probabilistic Processes. Theoretical Computer Science 538, pp. 16–36, doi:10.1016/j.tcs.2013.

07.016. Available at http://theory.stanford.edu/~rvg/abstracts.html#99.

[29] Y. Deng, R.J. van Glabbeek, M. Hennessy, C.C. Morgan & C. Zhang (2007): Characterising Testing Preorders for Finite Probabilistic Processes. In: Proceedings 22nd Annual IEEE Symposium on Logic in Computer Science, LICS’07, Wroclaw, Poland, July 2007, IEEE Computer Society

Press, pp. 313–322, doi:10.1109/LICS.2007.15. Available at http://theory.stanford.
edu/~rvg/abstracts.html#71.

[30] Y. Deng, R.J. van Glabbeek, M. Hennessy, C.C. Morgan & C. Zhang (2007): Remarks on Testing Probabilistic Processes. In L. Cardelli, M. Fiore & G. Winskel, editors: Computation, Meaning, and Logic: Articles dedicated to Gordon Plotkin, Electronic Notes in Theoretical Computer Science 172, Elsevier, pp. 359–397, doi:10.1016/j.entcs.2007.02.013. Available at http://

theory.stanford.edu/~rvg/abstracts.html#69.

[31] Y. Deng, R.J. van Glabbeek, C.C. Morgan & C. Zhang (2007): Scalar Outcomes Suffice for Finitary Probabilistic Testing. In R. De Nicola, editor: Proceedings 16th European Symposium on Programming, ESOP’07, Braga, Portugal, 24 March - 1 April, 2007, LNCS 4421, Springer, pp.

363–378, doi:10.1007/978-3-540-71316-6_25. Available at http://theory.stanford.
edu/~rvg/abstracts.html#70.

[32] V. Dyseryn, R.J. van Glabbeek & P. Höfner (2017): Analysing Mutual Exclusion using Process Algebra with Signals. In K. Peters & S. Tini, editors: Proceedings Combined 24th International Workshop on Expressiveness in Concurrency and 14th Workshop on Structural Operational Semantics, Berlin, Germany, 4th September 2017, Electronic Proceedings in Theoretical Computer Science 255, Open Publishing Association, pp. 18–34, doi:10.4204/EPTCS.255.2.

[33] A. Fehnker, R.J. van Glabbeek, P. Höfner, A.K. McIver, M. Portmann & W.L. Tan (2011): Modelling and Analysis of AODV in UPPAAL. In: Proceedings 1st International Workshop on Rigorous Protocol Engineering, WRiPE’11. Available at http://theory.stanford.edu/~rvg/

abstracts.html#91.

[34] A. Fehnker, R.J. van Glabbeek, P. Höfner, A.K. McIver, M. Portmann & W.L. Tan (2012): Automated Analysis of AODV Using UPPAAL. In C. Flanagan & B. König, editors: Proceedings 18th International Conference on Tools and Algorithms for the Construction and Analysis of Systems, TACAS’12; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’12, Tallinn, Estonia, March/April 2012, LNCS 7214, Springer, pp. 173–187,

doi:10.1007/978-3-642-28756-5_13. Available at http://theory.stanford.edu/~rvg/
abstracts.html#94.

[35] A. Fehnker, R.J. van Glabbeek, P. Höfner, A.K. McIver, M. Portmann & W.L. Tan (2012): A Process Algebra for Wireless Mesh Networks. In H. Seidl, editor: Programming Languages and Systems: Proceedings 21st European Symposium on Programming, ESOP’12; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’12, Tallinn, Estonia, March/April 2012, LNCS 7211, Springer, pp. 295–315, doi:10.1007/978-3-642-28869-2_15.

Available at http://theory.stanford.edu/~rvg/abstracts.html#93.

[36] A. Fehnker, R.J. van Glabbeek, P. Höfner, A.K. McIver, M. Portmann & W.L. Tan (2013): A Process Algebra for Wireless Mesh Networks used for Modelling, Verifying and Analysing AODV. Technical Report 5513, NICTA, Sydney, Australia. Available at http://arxiv.org/abs/1312.

7645.

[37] N. Fischer & R.J. van Glabbeek (2019): Axiomatising Infinitary ProbabilisticWeak Bisimilarity of Finite-State Behaviours. Journal of Logical and Algebraic Methods in Programming 102, pp.

64–102, doi:10.1016/j.jlamp.2018.09.006. Available at https://arxiv.org/abs/1810.
07488.

[38] W.J. Fokkink & R.J. van Glabbeek (1996): Ntyft/ntyxt rules reduce to ntree rules. Information and Computation 126(1), pp. 1–10, doi:10.1006/inco.1996.0030. Available at http://theory.

stanford.edu/~rvg/abstracts.html#33.

[39] W.J. Fokkink & R.J. van Glabbeek (2016): Divide and Congruence II: Delay and Weak Bisimilarity. In: Proceedings 31^{st}Annual ACM/IEEE Symposium on Logic in Computer Science, LICS’16, New York, NY, USA, July 2016, ACM, pp. 778–787, doi:10.1145/2933575.2933590. Avail-

able at http://theory.stanford.edu/~rvg/abstracts.html#116.

[40] W.J. Fokkink & R.J. van Glabbeek (2017): Divide and congruence II: From decomposition of modal formulas to preservation of delay and weak bisimilarity. Information and Computation

257, pp. 79–113, doi:10.1016/j.ic.2017.10.003. Available at http://arxiv.org/abs/
1604.07530.

[41] W.J. Fokkink & R.J. van Glabbeek (2017): Precongruence Formats with Lookahead through Modal Decomposition. In V. Goranko & M. Dam, editors: Proceedings 26th EACSL Annual Conference on Computer Science Logic, CSL’17, Leibniz International Proceedings in Informatics (LIPIcs) 82, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, Dagstuhl, Germany, pp.

25:1–25:20, doi:10.4230/LIPIcs.CSL.2017.25. Available at http://theory.stanford.
edu/~rvg/abstracts.html#122.

[42] W.J. Fokkink & R.J. van Glabbeek (2019): Preface. In W.J. Fokkink & R.J. van Glabbeek, editors: Proceedings 30th International Conference on Concurrency Theory (CONCUR 2019), Leibniz International Proceedings in Informatics (LIPIcs) 140, Schloss Dagstuhl–Leibniz-Zentrum fuer Informatik, Dagstuhl, Germany, pp. 0:i–0:xiv, doi:10.4230/LIPIcs.CONCUR.2019.0.

[43] W.J. Fokkink & R.J. van Glabbeek, editors (2019): Proceedings 30th International Conference on Concurrency Theory (CONCUR 2019). Leibniz International Proceedings in Informatics (LIPIcs) 140, Schloss Dagstuhl–Leibniz-Zentrum fuer Informatik, Dagstuhl, Germany. Available

at http://www.dagstuhl.de/dagpub/978-3-95977-121-4.

[44] W.J. Fokkink & R.J. van Glabbeek, editors (2021): Selected Papers of the 30th International Conference on Concurrency Theory (CONCUR 2019). Available at https://lmcs.episciences.

org/volume/view/id/386.

[45] W.J. Fokkink, R.J. van Glabbeek & B. Luttik (2017): Divide and Congruence III: From Decomposition of Modal Formulas to Preservation of Stability and Divergence. Technical Report, Data61,

CSIRO. Available at http://theory.stanford.edu/~rvg/abstracts.html#125.

[46] W.J. Fokkink, R.J. van Glabbeek & B. Luttik (2017): Divide and Congruence III: Stability & Divergence. In R. Meyer & U. Nestmann, editors: Proceedings 28th International Conference on Concurrency Theory, CONCUR’17, Leibniz International Proceedings in Informatics (LIPIcs) 85, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, Dagstuhl, Germany, pp. 15:1–

15:16, doi:10.4230/LIPIcs.CONCUR.2017.15. Available at http://theory.stanford.
edu/~rvg/abstracts.html#124.

[47] W.J. Fokkink, R.J. van Glabbeek & B. Luttik (2019): Divide and Congruence III: From Decomposition of Modal Formulas to Preservation of Stability and Divergence. Information and Computa-

tion 268:104435, doi:10.1016/j.ic.2019.104435. Available at https://arxiv.org/abs/
1908.06550.

[48] W.J. Fokkink, R.J. van Glabbeek & P. de Wind (2003): Compositionality of Hennessy-Milner Logic through Structural Operational Semantics. In A. Lingas & B.J. Nilsson, editors: Proceedings 14th International Symposium on Fundamentals of Computation Theory, FCT’03, Malmö, Sweden, August 2003, LNCS 2751, Springer, Berlin / Heidelberg, pp. 412–422, doi:10.1007/

978-3-540-45077-1_38. Available at http://theory.stanford.edu/~rvg/abstracts.
html#51.

[49] W.J. Fokkink, R.J. van Glabbeek & P. de Wind (2006): Compositionality of Hennessy-Milner Logic by Structural Operational Semantics. Theoretical Computer Science 354(3), pp. 421–

440, doi:10.1016/j.tcs.2005.11.035. Available at http://theory.stanford.edu/~rvg/
abstracts.html#61.

[50] W.J. Fokkink, R.J. van Glabbeek & P. de Wind (2006): Divide and Congruence Applied to η-Bisimulation. In P.D. Mosses & I. Ulidowski, editors: Proceedings of the Second Workshop on Structural Operational Semantics, SOS’05, Lisbon, Portugal, Electronic Notes in Theoretical Computer Science 156(1), Elsevier, pp. 97–113, doi:10.1016/j.entcs.2005.10.029. Avail-

able at http://theory.stanford.edu/~rvg/abstracts.html#62.

[51] W.J. Fokkink, R.J. van Glabbeek & P. de Wind (2006): Divide and Congruence: From Decomposition of Modalities to Preservation of Branching Bisimulation. In F.S. de Boer, M.M. Bonsangue, S. Graf & W.-P. de Roever, editors: Revised Lectures Fourth International Symposium on Formal Methods for Components and Objects, FMCO ’05, Amsterdam, The Netherlands, November 2005, LNCS 4111, Springer, pp. 195–218, doi:10.1007/11804192. Available at http://

theory.stanford.edu/~rvg/abstracts.html#64.

[52] W.J. Fokkink, R.J. van Glabbeek & P. de Wind (2012): Divide and congruence: From decomposition of modal formulas to preservation of branching and η-bisimilarity. Information and Computation 214, pp. 59–85, doi:10.1016/j.ic.2011.10.011. Available at http://theory.

stanford.edu/~rvg/abstracts.html#92.

[53] J. Gallagher, R.J. van Glabbeek & W. Serwe, editors (2018): Proceedings Third Workshop on Models for Formal Analysis of Real Systems and Sixth International Workshop on Verification and Program Transformation, Thessaloniki, Greece, 20th April 2018. Electronic Proceedings in Theoretical Computer Science 268, Open Publishing Association, doi:10.4204/EPTCS.268.

[54] R.J. van Glabbeek (2023): Reactive Bisimulation Semantics for a Process Algebra with Time-Outs. Acta Informatica 60, pp. 11–57, doi:10.1007/s00236-022-00417-1.

[55] R.J. van Glabbeek (1985): Good Coverings. Report nr. 3, Mathematical Institute, University of

Leiden, The Netherlands. Available at http://kilby.stanford.edu/~rvg/pub/good.pdf.

[56] R.J. van Glabbeek (1986): Bounded nondeterminism and the approximation induction principle in process algebra. Report CS-R8634, CWI, Amsterdam. Extended abstract in F.J. Brandenburg, G. Vidal-Naquet & M. Wirsing, editors: Proceedings STACS 87, 4^{th} Annual Symposium on Theoretical Aspects of Computer Science, Passau, Germany, February 1987, LNCS 247, Springer, 1987,

pp. 336–347, doi: 10.1007/BFb0039617.

[57] R.J. van Glabbeek (1987): Bounded nondeterminism and the approximation induction principle in process algebra (extended abstract). In F.J. Brandenburg, G. Vidal-Naquet & M. Wirsing, editors: Proceedings STACS 87, 4^{th} Annual Symposium on Theoretical Aspects of Computer Science, Passau, Germany, February 1987, LNCS 247, Springer, pp. 336–347, doi:10.1007/BFb0039617.

[58] R.J. van Glabbeek (1988): De semantiek van eindige, sequentiële processen met interne acties. Syllabus processemantieken, deel 2. Handwritten manuscript, in Dutch.

[59] R.J. van Glabbeek (1988): An operational non-interleaved process graph semantics of CCSP (abstract). In E.-R. Olderog, U. Goltz & R.J. van Glabbeek, editors: Combining Compositionality and Concurrency, Summary of a GMD-Workshop, Königswinter, March 1988, Arbeitspapiere der GMD 320, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin, pp. 18–19.

[60] R.J. van Glabbeek (1990): Comparative Concurrency Semantics and Refinement of Actions. Ph.D. thesis, Free University, Amsterdam. Introduction available at http://theory.stanford.edu/ ~rvg/thesis.html. Second edition available as CWI tract 109, CWI, Amsterdam 1996.

[61] R.J. van Glabbeek (1990): The Linear Time – Branching Time Spectrum. Report CS-R9029, CWI, Amsterdam. Extended abstract in J.C.M. Baeten & J.W. Klop, editors: Proceedings CONCUR ’90, Theories of Concurrency: Unification and Extension, Amsterdam, August 1990, LNCS 458, Springer, 1990, pp. 278–297, doi: 10.1007/BFb0039066.

[62] R.J. van Glabbeek (1990): The Linear Time – Branching Time Spectrum (extended abstract). In J.C.M. Baeten & J.W. Klop, editors: Proceedings CONCUR ’90, Theories of Concurrency: Unification and Extension, Amsterdam, August 1990, LNCS 458, Springer, pp. 278–297, doi:10.

1007/BFb0039066.

[63] R.J. van Glabbeek (1990): The refinement theorem for ST-bisimulation semantics. In M. Broy & C.B. Jones, editors: Proceedings IFIP TC2 Working Conference on Programming Concepts and Methods, Sea of Gallilee, Israel, April 1990, North-Holland, pp. 27–52. Available at http://

kilby.stanford.edu/~rvg/pub/STbisimulation.pdf.

[64] R.J. van Glabbeek (1991): Bisimulations for higher dimensional automata. Email message, July

7, 1991. Available at http://theory.stanford.edu/~rvg/hda.

[65] R.J. van Glabbeek (1993): A complete axiomatization for branching bisimulation congruence of finite-state behaviours. In A.M. Borzyszkowski & S. Soko�owski, editors: Proceedings 18^{th} International Symposium on Mathematical Foundations of Computer Science, MFCS ’93, Gdansk, Poland, August/September 1993, LNCS 711, Springer, pp. 473–484, doi:10.1007/

3-540-57182-5_39. Available at http://theory.stanford.edu/~rvg/abstracts.html#
25.

[66] R.J. van Glabbeek (1993): Full Abstraction in Structural Operational Semantics (extended abstract). In M. Nivat, C. Rattray, T. Rus & G. Scollo, editors: Proceedings of the 3^{rd} International Conference on Algebraic Methodology and Software Technology, AMAST’93, Twente, The Netherlands, June l993, Workshops in Computing, Springer, pp. 75–82. Available at http://

theory.stanford.edu/~rvg/abstracts.html#28.

[67] R.J. van Glabbeek (1993): The Linear Time – Branching Time Spectrum II; The semantics of sequential systems with silent moves (extended abstract). In E. Best, editor: Proceedings CON-CUR’93, 4^{th} International Conference on Concurrency Theory, Hildesheim, Germany, August 1993, LNCS 715, Springer, pp. 66–81, doi:10.1007/3-540-57208-2_6.

[68] R.J. van Glabbeek (1993): The linear time – branching time spectrum II; the semantics of sequential systems with silent moves (preliminary version). Manuscript. Available at http://theory. stanford.edu/~rvg/abstracts.html#26. Extended abstract in E. Best, editor: Proceedings CONCUR’93, 4^{th} International Conference on Concurrency Theory, Hildesheim, Germany, August 1993, LNCS 715, Springer, pp. 66–81, doi: 10.1007/3-540-57208-2_6.

[69] R.J. van Glabbeek (1994): On the expressiveness of ACP (extended abstract). In A. Ponse, C. Verhoef & S.F.M. van Vlijmen, editors: Proceedings First Workshop on the Algebra of Communicating Processes, ACP’94, Utrecht, The Netherlands, May 1994, Workshops in Computing, Springer,

pp. 188–217, doi:10.1007/978-1-4471-2120-6_8. Available at http://theory.stanford.
edu/~rvg/abstracts.html#31.

[70] R.J. van Glabbeek (1994): What is branching time semantics and why to use it? In M. Nielsen, editor: The Concurrency Column, Bulletin of the EATCS 53, pp. 190–198. Available at http:// Theory.stanford.edu/~rvg/branching. Also in G. Paun, G. Rozenberg & A. Salomaa, editors: Current Trends in Theoretical Computer Science; Entering the 21st Century, World Scientific, 2001, pp. 469-479.

[71] R.J. van Glabbeek (1995): Branching Bisimulation as a Tool in the Analysis of Weak Bisimulation.

Available at http://theory.stanford.edu/~rvg/abstracts.html#36.

[72] R.J. van Glabbeek (1995): The Meaning of Negative Premises in Transition System Specifications II. Technical Report STAN-CS-TN-95-16, Stanford University. Available at http://theory. stanford.edu/~rvg/abstracts.html#32. Extended abstract in F. Meyer auf der Heide & B. Monien, editors: Proceedings 23^{th} International Colloquium on Automata, Languages and Programming, ICALP ’96, Paderborn, Germany, July 1996, LNCS 1099, Springer, pp. 502–513,

doi: 10.1007/3-540-61440-0_154.

[73] R.J. van Glabbeek (1996): Comparative Concurrency Semantics and Refinement of Actions. CWI Tract 109, CWI, Amsterdam. Second edition of dissertation. Introduction available at http://

theory.stanford.edu/~rvg/thesis.html.

[74] R.J. van Glabbeek (1996): History preserving process graphs. Draft. Available at http://

theory.stanford.edu/~rvg/abstracts.html#hppg.

[75] R.J. van Glabbeek (1996): The Meaning of Negative Premises in Transition System Specifications II (extended abstract). In F. Meyer auf der Heide & B. Monien, editors: Proceedings 23^{th} International Colloquium on Automata, Languages and Programming, ICALP ’96, Paderborn, Germany, July 1996, LNCS 1099, Springer, pp. 502–513, doi:10.1007/3-540-61440-0_154. Available

at http://theory.stanford.edu/~rvg/abstracts.html#32.

[76] R.J. van Glabbeek (1996): Petri Nets, Configuration Structures, Propositional Theories and History Preserving Process Graphs (abstract). In M. Droste, E.-R. Olderog, B. Steffen & G. Winskel, editors: Semantics of Concurrent Systems—Foundations and Applications, Dagstuhl-Seminar-Report 144, Internationales Begegnungs- und Forschungszentrum für Informatik Schloss Dagstuhl, Postfach 15 11 50, D-66041 Saarbrücken, Germany, pp. 14–15. Available at http://

theory.stanford.edu/~rvg/nets-structures-graphs.

[77] R.J. van Glabbeek (1997): Axiomatizing Flat Iteration. In A. Mazurkiewicz & J. Winkowski, editors: Proceedings CONCUR ’97, 8^{th} International Conference on Concurrency Theory, Warsaw, Poland, July 1997, LNCS 1243, Springer, pp. 228–242, doi:10.1007/3-540-63141-0_16.

Available at http://theory.stanford.edu/~rvg/abstracts.html#38.

[78] R.J. van Glabbeek (1997): Notes on the methodology of CCS and CSP. Theoretical Computer Science 177(2), pp. 329–349, doi:10.1016/S0304-3975(96)00251-4. Available at http://

theory.stanford.edu/~rvg/abstracts.html#1. Originally appeared as Report CS-R8624,

CWI, Amsterdam, 1986.

[79] R.J. van Glabbeek (1998): On the Relative Expressiveness of Petri Nets, Event Structures and Process Algebras (abstract). In H.-D. Ehrich, U. Goltz & J. Meseguer, editors: Information Systems as Reactive Systems, Dagstuhl-Seminar-Report 200, Internationales Begegnungs- und Forschungszentrum für Informatik Schloss Dagstuhl, Postfach 15 11 50, D-66041 Saarbrücken,

Germany, p. 12. Available at http://theory.stanford.edu/~rvg/nets-process-alg.

[80] R.J. van Glabbeek (1999): Do we count from 0 or from 1? The ordinal use of cardinal expressions.

Position paper. Available at http://kilby.stanford.edu/~rvg/ordinal.html.

[81] R.J. van Glabbeek (1999): Petri Nets, Configuration Structures and Higher Dimensional Automata. In J.C.M. Baeten & S. Mauw, editors: Proceedings CONCUR ’99, 10^{th} International Conference on Concurrency Theory, Eindhoven, The Netherlands, August 1999, LNCS 1664, Springer, pp. 21–27, doi:10.1007/3-540-48320-9_3. Available at http://theory.

stanford.edu/~rvg/abstracts.html#42.

[82] R.J. van Glabbeek (1999): The third millennium starts on January 1 of the year 2001. Position

paper. Available at http://kilby.stanford.edu/~rvg/millennium.html.

[83] R.J. van Glabbeek (2001): Decidability, an introduction to decidability theory without invoking Church’ thesis. Class handout. Available at http://kilby.stanford.edu/~rvg/154/

handouts/decidability.html.

[84] R.J. van Glabbeek (2001): The Linear Time – Branching Time Spectrum I; The Semantics of Concrete, Sequential Processes. In J.A. Bergstra, A. Ponse & S.A. Smolka, editors: Handbook of Process Algebra, chapter 1, Elsevier, pp. 3–99, doi:10.1016/B978-044482830-9/50019-9.

Available at http://theory.stanford.edu/~rvg/abstracts.html#43.

[85] R.J. van Glabbeek (2001): The undefinability of definability. Class handout. Available at http://

kilby.stanford.edu/~rvg/154/handouts/definability.html.

[86] R.J. van Glabbeek (2001): What is branching time semantics and why to use it? In G. Paun, G. Rozenberg & A. Salomaa, editors: Current Trends in Theoretical Computer Science; Entering the 21st Century, World Scientific, pp. 469–479. Available at http://Theory.stanford.edu/ ~rvg/branching. Also in M. Nielsen, editor: The Concurrency Column, Bulletin of the EATCS 53, 1994, pp. 190–198.

[87] R.J. van Glabbeek (2003): Liveness respecting semantics. In L. Aceto, Z. Ésik, W.J. Fokkink & A. Ingólfsdóttir, editors: Slide Reprints from the Workshop on Process Algebra: Open Problems and Future Directions, PA ’03, Bologna, Italy, July 2003, BRICS notes NS-03-3, Department of Computer Science, University of Aarhus, Denmark, pp. 59–63. Available at http://www.

brics.dk/NS/03/3/BRICS-NS-03-3.pdf.

[88] R.J. van Glabbeek (2004): The Meaning of Negative Premises in Transition System Specifications II. Journal of Logic and Algebraic Programming 60–61, pp. 229–258, doi:10.1016/j.jlap.

2004.03.007. Available at http://theory.stanford.edu/~rvg/abstracts.html#53.

[89] R.J. van Glabbeek (2005): A Characterisation of Weak Bisimulation Congruence. In A. Middeldorp, V. van Oostrom, F. van Raamsdonk & R. de Vrijer, editors: Processes, Terms and Cycles: Steps on the Road to Infinity: Essays Dedicated to Jan Willem Klop on the Occasion of His 60th Birthday, LNCS 3838, Springer, pp. 26–39, doi:10.1007/11601548_4. Available at http://

theory.stanford.edu/~rvg/abstracts.html#65.

[90] R.J. van Glabbeek (2005): Higher-Dimensional Automata and Other Models of Concurrency. In P. Cousot, L. Fajstrup, E. Goubault, M. Herlihy, K.G. Larsen & M. Rauen, editors: Preliminary Proceedings of the Workshop on Geometry and Topology in Concurrency, GETCO ’05, San Francisco, USA, August 2005, BRICS Note NS-05-5, Department of Computer Science, University

of Aarhus, Denmark, p. 1. Available at http://theory.stanford.edu/~rvg/abstracts.
html#x.

[91] R.J. van Glabbeek (2005): The Individual and Collective Token Interpretations of Petri Nets. In M. Abadi & L. de Alfaro, editors: Proceedings CONCUR’05, 16^{th} International Conference on Concurrency Theory, San Francisco, USA, August 2005, LNCS 3653, Springer, pp.

323–337, doi:10.1007/11539452_26. Available at http://theory.stanford.edu/~rvg/
abstracts.html#59.

[92] R.J. van Glabbeek (2005): On Cool Congruence Formats for Weak Bisimulations (extended abstract). In D.V. Hung & M. Wirsing, editors: Proceedings International Colloquium on Theoretical Aspects of Computing, ICTAC’05, Hanoi, Vietnam, LNCS 3722, Springer, pp. 318–333, doi:10.

1007/11560647_21. Available at http://theory.stanford.edu/~rvg/abstracts.html#
58.

[93] R.J. van Glabbeek (2005): On Specifying Timeouts. In L. Aceto & A.D. Gordon, editors: Short Contributions from the Workshop on Algebraic Process Calculi: The First Twenty Five Years and Beyond, PA ’05, Bertinoro, Italy, August 2005, Electronic Notes in Theoretical Computer Science 162, Elsevier, pp. 112–113, doi:10.1016/j.entcs.2005.12.083. Available at http://

theory.stanford.edu/~rvg/abstracts.html#60.

[94] R.J. van Glabbeek (2005): On the Expressiveness of Higher Dimensional Automata (extended abstract). Electronic Notes in Theoretical Computer Science 128(2): Proceedings of the 11th International Workshop on Expressiveness in Concurrency (EXPRESS’04), pp. 5–34, doi:10.1016/

j.entcs.2004.11.026. Available at http://theory.stanford.edu/~rvg/abstracts.
html#56.

[95] R.J. van Glabbeek (2006): On the Expressiveness of Higher Dimensional Automata. Theoretical Computer Science 368(1-2), pp. 169–194, doi:10.1016/j.tcs.2006.06.024. Available at

http://theory.stanford.edu/~rvg/abstracts.html#63.

[96] R.J. van Glabbeek (2009): The Linear Time – Branching Time Spectrum after 20 years or Full abstraction for safety and liveness properties. Copies of slides. Invited talk for IFIP WG 1.8

at CONCUR’09 in Bologna. Available at http://theory.stanford.edu/~rvg/abstracts.
html#20years.

[97] R.J. van Glabbeek (2010): The Coarsest Precongruences Respecting Safety and Liveness Properties. In C.S. Calude & V. Sassone, editors: Proceedings 6th IFIP TC 1/WG 2.2 International Conference on Theoretical Computer Science (TCS’10); held as part of the World Computer Congress 2010, Brisbane, Australia, September 20-23, 2010, IFIP 323, Springer, pp. 32–52, doi:10.1007/

978-3-642-15240-5_3. Available at http://arxiv.org/abs/1007.5491.

[98] R.J. van Glabbeek (2011): Bisimulation. In D. Padua, editor: Encyclopedia of Parallel Computing, Springer, pp. 136–139, doi:10.1007/978-0-387-09766-4_149. Available at http://

theory.stanford.edu/~rvg/abstracts.html#45.

[99] R.J. van Glabbeek (2011): On Cool Congruence Formats for Weak Bisimulations. Theoretical Computer Science 412(28), pp. 3283–3302, doi:10.1016/j.tcs.2011.02.036. Available at

http://theory.stanford.edu/~rvg/abstracts.html#88.

[100] R.J. van Glabbeek (2012): Musings on Encodings and Expressiveness. In B. Luttik & M.A. Reniers, editors: Proceedings Combined 19th International Workshop on Expressiveness in Concurrency and 9th Workshop on Structured Operational Semantics, Newcastle upon Tyne, UK, September 3, 2012, Electronic Proceedings in Theoretical Computer Science 89, Open Publishing Association, pp. 81–98, doi:10.4204/EPTCS.89.7.

[101] R.J. van Glabbeek (2015): Structure Preserving Bisimilarity, Supporting an Operational Petri Net Semantics of CCSP. In R. Meyer, A. Platzer & H. Wehrheim, editors: Proceedings Correct System Design - Symposium in Honor of Ernst-Rüdiger Olderog on the Occasion of His 60th Birthday, Oldenburg, Germany, September 8-9, 2015, LNCS 9360, Springer, pp. 99–130, doi:10.1007/

978-3-319-23506-6_9. Available at http://arxiv.org/abs/1509.05842.

[102] R.J. van Glabbeek (2016): An Algebraic Treatment of Recursion. In I. Bethke, B. Bredeweg & A. Ponse, editors: Liber Amicorum for Jan A. Bergstra, Informatics Institute, University of

Amsterdam, pp. 58–59. Available at http://arxiv.org/abs/1702.07838.

[103] R.J. van Glabbeek (2016): Ensuring Liveness Properties of Distributed Systems (A Research Agenda). Position paper. Available at https://arxiv.org/abs/1711.04240.

[104] R.J. van Glabbeek (2017): A Branching Time Model of CSP. In Th. Gibson-Robinson, Ph.J. Hopcroft & R. Lazic, editors: Concurrency, Security, and Puzzles — Essays Dedicated to Andrew William Roscoe on the Occasion of His 60th Birthday, LNCS 10160, Springer, pp. 272–293,

doi:10.1007/978-3-319-51046-0_14. Available at http://arxiv.org/abs/1702.07844.

[105] R.J. van Glabbeek (2017): Lean and Full Congruence Formats for Recursion. In: Proceedings 32^{nd} Annual ACM/IEEE Symposium on Logic in Computer Science, LICS’17, Reykjavik, Iceland, June 2017, IEEE Computer Society Press, doi:10.1109/LICS.2017.8005142. Available

at https://arxiv.org/abs/1704.03160.

[106] R.J. van Glabbeek (2018): Is Speed-Independent Mutual Exclusion Implementable? In S. Schewe & L. Zhang, editors: Proceedings 29th International Conference on Concurrency Theory (CON-CUR’18), Beijing, China, September 2018, Leibniz International Proceedings in Informatics (LIPIcs) 118, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, Dagstuhl, Germany, doi:10.

4230/LIPIcs.CONCUR.2018.3.

[107] R.J. van Glabbeek (2018): On the Validity of Encodings of the Synchronous in the Asynchronous π-calculus. Information Processing Letters 137, p. 17–25, doi:10.1016/j.ipl.2018.04.015.

Available at https://arxiv.org/abs/1802.09182.

[108] R.J. van Glabbeek (2018): A Theory of Encodings and Expressiveness. Technical Report, Data61,

CSIRO. Available at https://arxiv.org/abs/1805.10415.

[109] R.J. van Glabbeek (2018): A Theory of Encodings and Expressiveness (extended abstract). In C. Baier & U. Dal Lago, editors: Proceeding 21st International Conference on Foundations of Software Science and Computational Structures, FoSSaCS’18; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’18, Thessaloniki, Greece, April 2018, LNCS 10803, Springer, pp. 183–202, doi:10.1007/978-3-319-89366-2_10.

[110] R.J. van Glabbeek (2019): Ensuring liveness properties of distributed systems: Open problems. Journal of Logical and Algebraic Methods in Programming 109:100480, doi:10.1016/j.jlamp.

2019.100480. Available at https://arxiv.org/abs/1912.05616.

[111] R.J. van Glabbeek (2019): Justness: A Completeness Criterion for Capturing Liveness Properties. Technical Report, Data61, CSIRO. Available at https://arxiv.org/abs/1909.00286. Extended abstract in M. Bojańczyk & A. Simpson, editors: Proceedings 22st International Conference on Foundations of Software Science and Computation Structures (FoSSaCS 2019); held as part of the European Joint Conferences on Theory and Practice of Software (ETAPS 2019), Prague, Czech Republic, April 2019, LNCS 11425, Springer, pp. 505-522.

[112] R.J. van Glabbeek (2019): Justness: A Completeness Criterion for Capturing Liveness Properties (extended abstract). In M. Bojańczyk & A. Simpson, editors: Proceedings 22st International Conference on Foundations of Software Science and Computation Structures (FoS-SaCS’19); held as part of the European Joint Conferences on Theory and Practice of Software (ETAPS’19), Prague, Czech Republic, April 2019, LNCS 11425, Springer, pp. 505–522, doi:10.

1007/978-3-030-17127-8_29.

[113] R.J. van Glabbeek (2019): On the Meaning of Transition System Specifications. In J.A. Pérez & J. Rot, editors: Proceedings Combined 26th International Workshop on Expressiveness in Concurrency and 16th Workshop on Structural Operational Semantics, Amsterdam, The Netherlands, 26th August 2019, Electronic Proceedings in Theoretical Computer Science 300, Open Publishing Association, pp. 69–85, doi:10.4204/EPTCS.300.5.

[114] R.J. van Glabbeek (2019): Reward Testing Equivalences for Processes. In M. Boreale, F. Corradini, M. Loreti & R. Pugliese, editors: Models, Languages, and Tools for Concurrent and Distributed Programming, Essays Dedicated to Rocco De Nicola on the Occasion of His 65th Birthday, LNCS 11665, Springer, pp. 45–70, doi:10.1007/978-3-030-21485-2_5. Available

at https://arxiv.org/abs/1907.13348.

[115] R.J. van Glabbeek (2020): Reactive Bisimulation Semantics for a Process Algebra with Time-Outs. In I. Konnov & L. Kovács, editors: Proceedings 31st International Conference on Concurrency Theory (CONCUR 20), Online, September 2020, Leibniz International Proceedings in Informatics (LIPIcs) 171, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, doi:10.4230/

LIPIcs.CONCUR.2020.6.

[116] R.J. van Glabbeek (2020): Reactive Temporal Logic. In O. Dardha & J. Rot, editors: Proceedings Combined 27th International Workshop on Expressiveness in Concurrency and 17th Workshop on Structural Operational Semantics, Online, 31 August 2020, Electronic Proceedings in Theoretical Computer Science 322, Open Publishing Association, pp. 51–68, doi:10.4204/EPTCS.322.6.

[117] R.J. van Glabbeek (2021): Coinductive Validity. Available at https://arxiv.org/abs/2104.

13021.

[118] R.J. van Glabbeek (2021): Comparing the expressiveness of the π-calculus and CCS. Available at http://arxiv.org/abs/2203.11519. Extended abstract in I. Sergey, editor: Programming Languages and Systems: Proceedings 31st European Symposium on Programming, ESOP’22; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’22, Munich, Germany, April 2022, LNCS 13240, Springer, 2022, pp. 548–574, doi: 10.

1007/978-3-030-99336-8_20.

[119] R.J. van Glabbeek (2021): Failure Trace Semantics for a Process Algebra with Time-outs. Logical Methods in Computer Science 17(2):11, doi:10.23638/LMCS-17(2:11)2021. Available at

http://arxiv.org/abs/2002.10814.

[120] R.J. van Glabbeek (2021): Modelling Mutual Exclusion in a Process Algebra with Time-outs.

arXiv:2106.12785v1.

[121] R.J. van Glabbeek (2021): Reactive Bisimulation Semantics for a Process Algebra with Time-Outs. Technical Report, Data61, CSIRO. Available at https://arxiv.org/abs/2008.11499. Extended abstract in I. Konnov & L. Kovács, editors: Proceedings 31st International Conference on Concurrency Theory (CONCUR 20), Online, September 2020, Leibniz International Proceedings in Informatics (LIPIcs) 171(6), Schloss Dagstuhl–Leibniz-Zentrum für Informatik.

[122] R.J. van Glabbeek (2022): Comparing the expressiveness of the π-calculus and CCS. In I. Sergey, editor: Programming Languages and Systems: Proceedings 31st European Symposium on Programming, ESOP’22; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’22, Munich, LNCS 13240, Springer, p. 548–574, doi:10.1007/

978-3-030-99336-8_20.

[123] R.J. van Glabbeek (2022): Fair Must Testing for I/O Automata. In N. Jansen, M. Stoelinga & P. van den Bos, editors: A Journey from Process Algebra via Timed Automata to Model Learning, LNCS 13560, Springer, pp. 559–574, doi:10.1007/978-3-031-15629-8_30. arXiv:2212.

11248.

[124] R.J. van Glabbeek (2023): Comparing the expressiveness of the π-calculus and CCS. ACM Transactions on Computational Logic, doi:10.1145/3611013.

[125] R.J. van Glabbeek (2023): Just Testing. In P. Sobocinski & O. Kupferman, editors: Proceedings 26th International Conference on Foundations of Software Science and Computation Structures (FoSSaCS 2023); held as part of the European Joint Conferences on Theory and Practice of Software (ETAPS 2023), Paris, France, April 2023, LNCS 13992, Springer, pp. 498–519, doi:10.

1007/978-3-031-30829-1_24. arXiv:2212.08829.

[126] R.J. van Glabbeek (2023): Modelling Mutual Exclusion in a Process Algebra with Time-outs. Information and Computation 294:105079, doi:10.1016/j.ic.2023.105079.

[127] R.J. van Glabbeek & U. Goltz (1989): Equivalence notions for concurrent systems and refinement of actions. Arbeitspapiere der GMD 366, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin. Extended abstract in A. Kreczmar & G. Mirkowska: Proceedings 14^{th} Symposium on Mathematical Foundations of Computer Science, MFCS ’89, Pora̧bka-Kozubnik, Poland, August/September 1989, LNCS 379, Springer, pp. 237–248, doi: 10.1007/3-540-51486-4_

71.

[128] R.J. van Glabbeek & U. Goltz (1989): Equivalence notions for concurrent systems and refinement of actions (extended abstract). In A. Kreczmar & G. Mirkowska, editors: Proceedings 14^{th} Symposium on Mathematical Foundations of Computer Science, MFCS ’89, Pora̧bka-Kozubnik, Poland, August/September 1989, LNCS 379, Springer, pp. 237–248, doi:10.1007/

3-540-51486-4_71.

[129] R.J. van Glabbeek & U. Goltz (1989): Partial order semantics for refinement of actions—neither necessary nor always sufficient but appropriate when used with care. Bulletin of the European Association for Theoretical Computer Science 38, pp. 154–163.

[130] R.J. van Glabbeek & U. Goltz (1990): A Deadlock-sensitive Congruence for Action Refinement. SFB-Bericht Nr. 342/23/90 A, Institut für Informatik, Technische Universität München. Abstract in E. Best & G. Rozenberg, editors: Proceedings 3^{rd} Workshop on Concurrency and Compositionality, Goslar, Germany, March 1991, GMD-Studien Nr. 191, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin 1991, pp. 113–116.

[131] R.J. van Glabbeek & U. Goltz (1990): Equivalences and Refinement. In I. Guessarian, editor: Proceedings Semantics of Systems of Concurrent Processes, LITP Spring School on Theoretical Computer Science, La Roche Posay, France, April 1990, LNCS 469, Springer, pp. 309–333,

doi:10.1007/3-540-53479-2_13.

[132] R.J. van Glabbeek & U. Goltz (1990): Refinement of actions in causality based models. In J.W. de Bakker, W.P. de Roever & G. Rozenberg, editors: Proceedings REX Workshop on Stepwise Refinement of Distributed Systems: Models, Formalism, Correctness, Mook, The Netherlands, May/June 1989, LNCS 430, Springer, pp. 267–300, doi:10.1007/3-540-52559-9_68.

[133] R.J. van Glabbeek & U. Goltz (1991): A Deadlock-sensitive Congruence for Action Refinement (abstract). In E. Best & G. Rozenberg, editors: Proceedings 3^{rd} Workshop on Concurrency and Compositionality, Goslar, Germany, March 1991, GMD-Studien Nr. 191, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin, pp. 113–116.

[134] R.J. van Glabbeek & U. Goltz (2001): Refinement of Actions and Equivalence Notions for Concurrent Systems. Acta Informatica 37, pp. 229–327, doi:10.1007/s002360000041. Available at

http://theory.stanford.edu/~rvg/abstracts.html#41.

[135] R.J. van Glabbeek & U. Goltz (2004): Well-behaved Flow Event Structures for Parallel Composition and Action Refinement. Theoretical Computer Science 311, pp. 463–478, doi:10.1016/j.

tcs.2003.10.031. Available at http://theory.stanford.edu/~rvg/abstracts.html#
46.

[136] R.J. van Glabbeek, U. Goltz, C. Lippert & S. Mennicke (2019): Stronger Validity Criteria for Encoding Synchrony. In M.S. Alvim, K. Chatzikokolakis, C. Olarte & F. Valencia, editors: The Art of Modelling Computational Systems: A Journey from Logic and Concurrency to Security and Privacy - Essays Dedicated to Catuscia Palamidessi on the Occasion of Her 60th Birthday, LNCS 11760, Springer, pp. 182–205, doi:10.1007/978-3-030-31175-9_11.

[137] R.J. van Glabbeek, U. Goltz & E.-R. Olderog (2015): Special issue on “Combining Compositionality and Concurrency”: part 1. Acta Informatica 52(1), pp. 3–4, doi:10.1007/

s00236-014-0213-y.

[138] R.J. van Glabbeek, U. Goltz & E.-R. Olderog (2015): Special issue on “Combining Compositionality and Concurrency”: part 2. Acta Informatica 52(4-5), pp. 303–304, doi:10.1007/

s00236-015-0240-3.

[139] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2008): On Synchronous and Asynchronous Interaction in Distributed Systems. Technical Report 2008-04, Technical University of Braunschweig. Available at http://arxiv.org/abs/0901.0048. Extended abstract in: E. Ochmański & J. Tyszkiewicz, editors: Proceedings 33rd International Symposium on Mathematical Foundations of Computer Science (MFCS’08), Toruń, Poland, August 2008, LNCS 5162, Springer, pp. 16–35.

[140] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2008): On Synchronous and Asynchronous Interaction in Distributed Systems. In E. Ochmański & J. Tyszkiewicz, editors: Proceedings 33rd International Symposium on Mathematical Foundations of Computer Science (MFCS’08), Toruń, Poland, August 2008, LNCS 5162, Springer, pp. 16–35, doi:10.1007/978-3-540-85238-4_2.

Available at http://theory.stanford.edu/~rvg/abstracts.html#78.

[141] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2008): Symmetric and Asymmetric Asynchronous Interaction. Technical Report 2008-03, Technical University of Braunschweig. Available at http://arxiv.org/abs/0901.0043. Extended abstract in: F. Bonchi, D. Grohmann, P. Spoletini, A. Troina & E. Tuosto, editors: Proceedings of the First Interaction and Concurrency Experience (ICE’08), ENTCS 229(3), pp. 77-95.

[142] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2009): Symmetric and Asymmetric Asynchronous Interaction. In F. Bonchi, D. Grohmann, P. Spoletini, A. Troina & E. Tuosto, editors: Proceedings of the First Interaction and Concurrency Experience (ICE’08), Electronic Notes in Theoretical Computer Science 229(3), Elsevier, pp. 77–95, doi:10.1016/j.entcs.2009.06.040. Available

at http://theory.stanford.edu/~rvg/abstracts.html#77.

[143] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2011): Abstract Processes of Place/Transition Systems. Information Processing Letters 111(13), pp. 626–633, doi:10.1016/j.ipl.2011.03.

013. Available at http://arxiv.org/abs/1103.5916.

[144] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2011): On Causal Semantics of Petri Nets. Informatik Bericht Nr. 2011-06, Institut für Programmierung und Reaktive Systeme, TU Braun-

schweig, Germany. Available at http://theory.stanford.edu/~rvg/abstracts.html#90.

Extended abstract in J.-P. Katoen and B. König, editors: Proceedings 22nd International Conference on Concurrency Theory, CONCUR’11, Aachen, Germany, September 2011, LNCS 6901, Springer, 2011, pp. 43-59.

[145] R.J. van Glabbeek, U. Goltz & J.-W. Schicke (2011): On Causal Semantics of Petri Nets (extended abstract). In J.-P. Katoen & B. König, editors: Proceedings 22nd International Conference on Concurrency Theory, CONCUR’11, Aachen, Germany, September 2011, LNCS 6901, Springer,

pp. 43–59, doi:10.1007/978-3-642-23217-6_4. Available at http://theory.stanford.
edu/~rvg/abstracts.html#90.

[146] R.J. van Glabbeek, U. Goltz & J.-W. Schicke-Uffmann (2012): On Distributability of Petri Nets. Informatik Bericht Nr. 2011-10, Institut für Programmierung und Reaktive Systeme, TU Braunschweig, Germany. Available at http://arxiv.org/abs/1207.3597. Extended abstract in L. Birkedal, editor: Proceeding 15th International Conference on Foundations of Software Science and Computational Structures, FoSSaCS’12; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’12, Tallinn, Estonia, March/April 2012, LNCS 7213, Springer, 2012, pp. 331-345.

[147] R.J. van Glabbeek, U. Goltz & J.-W. Schicke-Uffmann (2012): On Distributability of Petri Nets (extended abstract). In L. Birkedal, editor: Proceeding 15th International Conference on Foundations of Software Science and Computational Structures, FoSSaCS’12; held as part of the European Joint Conferences on Theory and Practice of Software, ETAPS’12, Tallinn, Estonia, March/April 2012, LNCS 7213, Springer, pp. 331–345, doi:10.1007/978-3-642-28729-9_22.

Available at http://theory.stanford.edu/~rvg/abstracts.html#95.

[148] R.J. van Glabbeek, U. Goltz & J.-W. Schicke-Uffmann (2013): On Characterising Distributability. Logical Methods in Computer Science 9(3):17, doi:10.2168/LMCS-9(3:17)2013. Available

at http://arxiv.org/abs/1309.3883.

[149] R.J. van Glabbeek, U. Goltz & J.-W. Schicke-Uffmann (2021): Abstract Processes and Conflicts in Place/Transition Systems. Information and Computation 281:104706, doi:10.1016/j.ic.2021.

104706. arXiv:2103.01490.

[150] R.J. van Glabbeek, U. Goltz & J.-W. Schicke-Uffmann (2022): Abstract Processes in the Absence of Conflicts in General Place/Transition Systems. Information and Computation 289A:104939,

doi:10.1016/j.ic.2022.104939. arXiv:2207.04362.

[151] R.J. van Glabbeek, V. Gramoli & P. Tholoniat (2019): Cross-Chain Payment Protocols with Suc-

cess Guarantees. Available at http://arxiv.org/abs/1912.04513.

[152] R.J. van Glabbeek, V. Gramoli & P. Tholoniat (2020): Feasibility of Cross-Chain Payment with Success Guarantees. In: Proceedings 32nd ACM Symposium on Parallelism in Algorithms and Architectures, SPAA 2020, Virtual Event, USA, July 2020, ACM, pp. 579–581, doi:10.1145/

3350755.3400264. Available at http://arxiv.org/abs/2007.08152.

[153] R.J. van Glabbeek, V. Gramoli & P. Tholoniat (2023): Cross-Chain Payment Protocols with Success Guarantees. Distributed Computing 26, p. 137–157, doi:10.1007/s00446-023-00446-0.

[154] R.J. van Glabbeek, J.F. Groote & P. Höfner (2015): Preface, Proceedings Workshop on Models for Formal Analysis of Real Systems, Suva, Fiji, November 23, 2015. Electronic Proceedings in Theoretical Computer Science 196:0, doi:10.4204/EPTCS.196.0.

[155] R.J. van Glabbeek, J.F. Groote & P. Höfner, editors (2015): Proceedings Workshop on Models for Formal Analysis of Real Systems, Suva, Fiji, November 23, 2015. Electronic Proceedings in Theoretical Computer Science 196, Open Publishing Association, doi:10.4204/EPTCS.196.

[156] R.J. van Glabbeek, J.F. Groote & E.P. de Vink (2019): A Complete Axiomatization of Branching Bisimilarity for a Simple Process Language with Probabilistic Choice (Extended Abstract). In M.S. Alvim, K. Chatzikokolakis, C. Olarte & F. Valencia, editors: The Art of Modelling Computational Systems: A Journey from Logic and Concurrency to Security and Privacy - Essays Dedicated to Catuscia Palamidessi on the Occasion of Her 60th Birthday, LNCS 11760, Springer,

pp. 139–162, doi:10.1007/978-3-030-31175-9_9.

[157] R.J. van Glabbeek, J.F. Groote & E.P. de Vink (2023): A Cancellation Law for Probabilistic Processes. In C.A. Mezzina & G. Caltais, editors: Proceedings Combined 30th International Workshop on Expressiveness in Concurrency and 20th Workshop on Structural Operational Semantics, Antwerp, Belgium, 18th September 2023, Electronic Proceedings in Theoretical Computer Science 387, Open Publishing Association, pp. 42–58, doi:10.4204/EPTCS.387.5.

[158] R.J. van Glabbeek & M. Hennessy, editors (2006): Proceedings of the 4th Workshop on Structural Operational Semantics, Wroclaw, Poland, 9 July 2007. Electronic Notes in Theoretical Computer

Science 192(1), Elsevier. Available at http://theory.stanford.edu/~rvg/abstracts.
html#72.

[159] R.J. van Glabbeek & M. Hennessy (2007): Preface Proceedings of the 4th Workshop on Structural Operational Semantics, Wroclaw, Poland, 9 July 2007. Electronic Notes in Theoretical Computer

Science 192(1), pp. 1–3, doi:10.1016/j.entcs.2007.10.001. Available at http://theory.
stanford.edu/~rvg/abstracts.html#72.

[160] R.J. van Glabbeek & P. Höfner (2015): CCS: It’s not fair! - Fair schedulers cannot be implemented in CCS-like languages even under progress and certain fairness assumptions. Acta Informatica

52(2-3), pp. 175–205, doi:10.1007/s00236-015-0221-6. Available at http://arxiv.org/
abs/1505.05964.

[161] R.J. van Glabbeek & P. Höfner (2015): Progress, Fairness and Justness in Process Algebra. Technical Report 8501, NICTA, Sydney, Australia. Available at http://arxiv.org/abs/1501.

03268.

[162] R.J. van Glabbeek & P. Höfner (2017): Split, Send, Reassemble: A Formal Specification of a CAN Bus Protocol Stack. In H. Hermanns & P. Höfner, editors: Proceedings 2nd Workshop on Models for Formal Analysis of Real Systems, Uppsala, Sweden, 29th April 2017, Electronic Proceedings in Theoretical Computer Science 244, Open Publishing Association, pp. 14–52, doi:10.4204/

EPTCS.244.2.

[163] R.J. van Glabbeek & P. Höfner (2018): Progress, Justness and Fairness. Survey paper, Data61, CSIRO, Sydney, Australia. Available at https://arxiv.org/abs/1810.07414v1.

[164] R.J. van Glabbeek & P. Höfner (2019): Progress, Justness and Fairness. ACM Computing Surveys

52(4):69, doi:10.1145/3329125. Available at https://arxiv.org/abs/1810.07414.

[165] R.J. van Glabbeek, P. Höfner & R. Horne (2021): Assuming Just Enough Fairness to make Session Types Complete for Lock-freedom. In: Proceedings 36^{th}Annual ACM/IEEE Symposium on Logic in Computer Science, LICS’21, IEEE Computer Society Press, doi:10.1109/LICS52264.2021.

9470531. Available at https://arxiv.org/abs/2104.14226.

[166] R.J. van Glabbeek, P. Höfner & M. Markl (2019): A Process Algebra for Link Layer Protocols. In L. Caires, editor: Programming Languages and Systems, Proceedings 28th European Symposium on Programming (ESOP’19); held as part of the European Joint Conferences on Theory and Practice of Software (ETAPS’19), Prague, Czech Republic, April 2019, LNCS 11423, Springer,

pp. 668–693, doi:10.1007/978-3-030-17184-1_24. Available at https://arxiv.org/abs/
1907.13329.

[167] R.J. van Glabbeek, P. Höfner, M. Portmann & W.L. Tan (2016): Modelling and Verifying the AODV Routing Protocol. Distributed Computing 29(4), pp. 279–315, doi:10.1007/

s00446-015-0262-7. Available at http://arxiv.org/abs/1512.08867.

[168] R.J. van Glabbeek, P. Höfner, W.L. Tan & M. Portmann (2013): Sequence Numbers Do Not Guarantee Loop Freedom —AODV Can Yield Routing Loops—. In: Proceedings 16th ACM International Conference on Modeling, Analysis and Simulation of Wireless and Mobile Systems, MSWiM ’13, Barcelona, Spain, November 2013, ACM, pp. 91–100, doi:10.1145/2507924.

2507943. Available at http://theory.stanford.edu/~rvg/abstracts.html#100.

[169] R.J. van Glabbeek, P. Höfner & D. van der Wal (2018): Analysing AWN-Specifications Using mCRL2 (Extended Abstract). In C.A. Furia & K. Winter, editors: Proceedings 14th International Conference on Integrated Formal Methods (IFM’18), Maynooth, Ireland, September 2018, LNCS 11023, Springer, pp. 398–418, doi:10.1007/978-3-319-98938-9_23. Available at http://

theory.stanford.edu/~rvg/abstracts.html#132.

[170] R.J. van Glabbeek, P. Höfner & W. Wang (2021): Enabling Preserving Bisimulation Equivalence. In S. Haddad & D. Varacca, editors: Proceedings 32nd International Conference on Concurrency Theory, CONCUR’21, Leibniz International Proceedings in Informatics (LIPIcs) 203, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, doi:10.4230/LIPIcs.CONCUR.2021.33.

[171] R.J. van Glabbeek, P. Höfner & W. Wang (2021): Enabling Preserving Bisimulation Equivalence. Available at https://arxiv.org/abs/2108.00142. Extended abstract in: Proceedings 32nd International Conference on Concurrency Theory, CONCUR’21, Leibniz International Proceedings in Informatics (LIPIcs) 203, Schloss Dagstuhl–Leibniz-Zentrum für Informatik, 2021,

doi:10.4230/LIPIcs.CONCUR.2021.33.

[172] R.J. van Glabbeek, P. Höfner & W. Wang (2023): A Lean-Congruence Format for EP-Bisimilarity. In C.A. Mezzina & G. Caltais, editors: Proceedings Combined 30th International Workshop on Expressiveness in Concurrency and 20th Workshop on Structural Operational Semantics, Antwerp, Belgium, 18th September 2023, Electronic Proceedings in Theoretical Computer Science 387, Open Publishing Association, pp. 59–75, doi:10.4204/EPTCS.387.6.

[173] R.J. van Glabbeek & D.J.D. Hughes (2016): MALL proof nets identify proofs modulo rule com-

mutation. Available at http://theory.stanford.edu/~rvg/abstracts.html#118.

[174] R.J. van Glabbeek, B. Luttik & L. Spaninks (2018): Rooted Divergence-Preserving Branching Bisimilarity is a Congruence. Technical Report, Data61, CSIRO. Available at https://arxiv.

org/abs/1801.01180v1.

[175] R.J. van Glabbeek, B. Luttik & L. Spaninks (2020): Rooted Divergence-Preserving Branching Bisimilarity is a Congruence. Logical Methods in Computer Science 16(3):14, doi:10.23638/

LMCS-16(3:14)2020. Available at https://arxiv.org/abs/1801.01180.

[176] R.J. van Glabbeek, B. Luttik & N. Trčka (2009): Branching Bisimilarity with Explicit Divergence. Fundamenta Informaticae 93(4), pp. 371–392, doi:10.3233/FI-2009-109. Available at

http://theory.stanford.edu/~rvg/abstracts.html#67.

[177] R.J. van Glabbeek, B. Luttik & N. Trčka (2009): Computation Tree Logic with Deadlock Detection. Logical Methods in Computer Science 5(4), doi:10.2168/LMCS-5(4:5)2009. Available at

http://theory.stanford.edu/~rvg/abstracts.html#73.

[178] R.J. van Glabbeek & C.A. Middelburg (2020): On Infinite Guarded Recursive Specifications in

Process Algebra. Available at http://arxiv.org/abs/2005.00746.

[179] R.J. van Glabbeek & P.D. Mosses (2006): Preface Proceedings of the 3rd Workshop on Structural Operational Semantics, Bonn, Germany, 26 August 2006. Electronic Notes in Theoretical Computer Science 175(1), pp. 1–2, doi:10.1016/j.entcs.2006.11.016. Available at http://

theory.stanford.edu/~rvg/abstracts.html#68.

[180] R.J. van Glabbeek & P.D. Mosses, editors (2006): Proceedings of the 3rd Workshop on Structural Operational Semantics, Bonn, Germany, 26 August 2006. Electronic Notes in Theoretical Computer Science 175(1), Elsevier. Available at http://theory.stanford.edu/~rvg/

abstracts.html#68.

[181] R.J. van Glabbeek & P.D. Mosses (2009): Preface, Special Issue on Structural Operational Semantics. Information and Computation 207(2), pp. 83–84, doi:10.1016/j.ic.2008.10.006.

Available at http://theory.stanford.edu/~rvg/abstracts.html#76.

[182] R.J. van Glabbeek & B. Ploeger (2008): Correcting a Space-Efficient Simulation Algorithm. In A. Gupta & S. Malik, editors: Proceedings 20th International Conference on Computer Aided Verification (CAV’08), Princeton, USA, July 2008, LNCS 5123, Springer, pp. 517–529,

doi:10.1007/978-3-540-70545-1_49. Available at http://theory.stanford.edu/~rvg/
abstracts.html#74.

[183] R.J. van Glabbeek & B. Ploeger (2008): Correcting a Space-Efficient Simulation Algorithm. CS-Report 08-06, Eindhoven University of Technology. Available at http://theory.stanford. edu/~rvg/abstracts.html#74. Extended abstract in A. Gupta & S. Malik, editors: Proceedings 20th International Conference on Computer Aided Verification (CAV’08), Princeton, USA, July 2008, LNCS 5123, Springer, pp. 517–529, doi: 10.1007/978-3-540-70545-1_49.

[184] R.J. van Glabbeek & B. Ploeger (2008): Five Determinisation Algorithms. In O.H. Ibarra & B. Ravikumar, editors: Proceedings Thirteenth International Conference on Implementation and Application of Automata (CIAA’08), San Francisco, California, USA, July 2008, LNCS 5148,

Springer, pp. 161–170, doi:10.1007/978-3-540-70844-5_17. Available at http://theory.
stanford.edu/~rvg/abstracts.html#75.

[185] R.J. van Glabbeek & B. Ploeger (2008): Five Determinisation Algorithms. CS-Report 08-14, Eindhoven University of Technology. Available at http://theory.stanford.edu/~rvg/ abstracts.html#75. Extended abstract in A. Gupta & S. Malik, editors: Proceedings Thirteenth International Conference on Implementation and Application of Automata (CIAA’08), San Francisco, California, USA, July 2008, LNCS 5148, Springer, pp. 161–170, doi: 10.1007/

978-3-540-70844-5_17.

[186] R.J. van Glabbeek & G.D. Plotkin (1995): Configuration Structures (extended abstract). In D. Kozen, editor: Proceedings 10^{th} Annual IEEE Symposium on Logic in Computer Science, LICS’95, San Diego, USA, June 1995, IEEE Computer Society Press, pp. 199–209, doi:10.1109/

LICS.1995.523257. Available at http://theory.stanford.edu/~rvg/abstracts.html#
34.

[187] R.J. van Glabbeek & G.D. Plotkin (2004): Event Structures for Resolvable Conflict. In J. Fiala, V. Koubek & J. Kratochvíl, editors: Proceedings 29^{th}International Symposium on Mathematical Foundations of Computer Science, MFCS’04, Prague, Czech Republic, August 2004, LNCS 3153,

Springer, pp. 550–561, doi:10.1007/978-3-540-28629-5_42. Available at http://theory.
stanford.edu/~rvg/abstracts.html#55.

[188] R.J. van Glabbeek & G.D. Plotkin (2009): Configuration structures, event structures and Petri nets. Theoretical Computer Science 410(41), pp. 4111–4159, doi:10.1016/j.tcs.2009.06.

014. Available at http://arxiv.org/abs/0912.4023.

[189] R.J. van Glabbeek & G.D. Plotkin (2010): On CSP and the Algebraic Theory of Effects. In C.B. Jones, A.W. Roscoe & K.R. Wood, editors: Reflections on the Work of C.A.R. Hoare, History of Computing, Springer, pp. 333–369, doi:10.1007/978-1-84882-912-1_15. Available at

http://arxiv.org/abs/1007.5488.

[190] R.J. van Glabbeek & P. Rittgen (1998): Scheduling Algebra. Arbeitsberichte des Instituts für Wirtschaftsinformatik 12, Universität Koblenz-Landau, Germany. Available at http://theory. stanford.edu/~rvg/abstracts.html#40. Slightly revised version in A.M. Haeberer, editor: Proceedings of the Seventh International Conference on Algebraic Methodology and Software Technology, AMAST’98, Amazonia, Brazil, January 1999, LNCS 1548, Springer, 1999, pp. 278–

292, doi: 10.1007/3-540-49253-4_21.

[191] R.J. van Glabbeek & P. Rittgen (1998): Scheduling Algebra. Technical Report STAN-CS-TN-

98-87, Stanford University. Available at http://theory.stanford.edu/~rvg/abstracts.

html#40. Slightly condensed version in A.M. Haeberer, editor: Proceedings of the Seventh International Conference on Algebraic Methodology and Software Technology, AMAST’98, Amazonia, Brazil, January 1999, LNCS 1548, Springer, 1999, pp. 278–292, doi: 10.1007/

3-540-49253-4_21.

[192] R.J. van Glabbeek & P. Rittgen (1999): Scheduling Algebra. In A.M. Haeberer, editor: Proceedings of the 7^{th} International Conference on Algebraic Methodology and Software Technology, AMAST’98, Amazonia, Brazil, January 1999, LNCS 1548, Springer, pp. 278–292, doi:10.1007/

3-540-49253-4_21. Available at http://theory.stanford.edu/~rvg/abstracts.html#
40.

[193] R.J. van Glabbeek & J.J.M.M. Rutten (1989): The processes of De Bakker and Zucker represent bisimulation equivalence classes. In: J.W. de Bakker, 25 jaar semantiek, liber amicorum, CWI, Amsterdam, pp. 243–246.

[194] R.J. van Glabbeek, S.A. Smolka & B. Steffen (1995): Reactive, generative, and stratified models of probabilistic processes. Information and Computation 121(1), pp. 59–80, doi:10.1006/inco.

1995.1123. Available at http://theory.stanford.edu/~rvg/abstracts.html#30.

[195] R.J. van Glabbeek, S.A. Smolka, B. Steffen & C.M.N. Tofts (1990): Reactive, generative, and stratified models of probabilistic processes. In: Proceedings 5^{th} Annual IEEE Symposium on Logic in Computer Science, LICS’90, Philadelphia, USA, June 1990, IEEE Computer Society

Press, pp. 130–141, doi:10.1109/LICS.1990.113740.

[196] R.J. van Glabbeek & D.G. Stork (2003): Query Nets: Interacting Workflow Modules that Ensure Global Termination. In W.M.P. van der Aalst, A.H.M. ter Hofstede & M. Weske, editors: Proceedings International Conference on Business Process Management, BPM’03, Eindhoven, The Netherlands, June 2003, LNCS 2678, Springer, pp. 184–199, doi:10.1007/3-540-44895-0_13.

Available at http://theory.stanford.edu/~rvg/abstracts.html#49.

[197] R.J. van Glabbeek & F.W. Vaandrager (1987): Petri net models for algebraic theories of concurrency (extended abstract). In J.W. de Bakker, A.J. Nijman & P.C. Treleaven, editors: Proceedings PARLE, Parallel Architectures and Languages Europe, Eindhoven, The Netherlands, June 1987, Vol. II: Parallel Languages, LNCS 259, Springer, pp. 224–242, doi:10.1007/3-540-17945-3_

13. Available at http://kilby.stanford.edu/~rvg/pub/petri.pdf.

[198] R.J. van Glabbeek & F.W. Vaandrager (1988): Modular specifications in process algebra—with curious queues. Report CS-R8821, CWI, Amsterdam. Extended abstract in M. Wirsing & J.A. Bergstra, editors: Algebraic Methods: Theory, Tools and Applications, LNCS 394, Springer, 1989, pp. 465–506, doi: 10.1007/BFb0015049.

[199] R.J. van Glabbeek & F.W. Vaandrager (1989): Modular specifications in process algebra—with curious queues (extended abstract). In M. Wirsing & J.A. Bergstra, editors: Algebraic Methods: Theory, Tools and Applications, LNCS 394, Springer, pp. 465–506, doi:10.1007/BFb0015049.

[200] R.J. van Glabbeek & F.W. Vaandrager (1991): The Difference Between Splitting in n and n+1 (abstract). In E. Best & G. Rozenberg, editors: Proceedings 3^{rd} Workshop on Concurrency and Compositionality, Goslar, Germany, March 1991, GMD-Studien Nr. 191, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin, pp. 117–121.

[201] R.J. van Glabbeek & F.W. Vaandrager (1993): Modular Specification of Process Algebras. Theoretical Computer Science 113(2), pp. 293–348, doi:10.1016/0304-3975(93)90006-F.

[202] R.J. van Glabbeek & F.W. Vaandrager (1997): The Difference Between Splitting in n and n+1. Information and Computation 136(2), pp. 109–142, doi:10.1006/inco.1997.2634. Available

at http://boole.stanford.edu/pub/split.pdf. Abstract in E. Best & G. Rozenberg, edi-

tors: Proceedings 3^{rd} Workshop on Concurrency and Compositionality, Goslar, Germany, March 1991, GMD-Studien Nr. 191, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin 1991, pp. 117–121.

[203] R.J. van Glabbeek & F.W. Vaandrager (2003): Bundle Event Structures and CCSP. In R. Amadio & D. Lugiez, editors: Proceedings CONCUR’03, 14^{th} International Conference on Concurrency Theory, Marseille, France, September 2003, LNCS 2761, Springer, pp. 57–71, doi:10.1007/

978-3-540-45187-7_4. Available at http://theory.stanford.edu/~rvg/abstracts.
html#52.

[204] R.J. van Glabbeek & M. Voorhoeve (2006): Liveness, Fairness and Impossible Futures. In C. Baier & H. Hermanns, editors: Proceedings CONCUR’06, 17^{th} International Conference on Concurrency Theory, Bonn, Germany, August 2006, LNCS 4137, Springer, pp. 126–141, doi:10.

1007/11817949_9. Available at http://theory.stanford.edu/~rvg/abstracts.html#
66.

[205] R.J. van Glabbeek & W.P. Weijland (1989): Branching Time and Abstraction in Bisimulation Semantics (extended abstract). In G.X. Ritter, editor: Information Processing 89, Proceedings of the IFIP 11th World Computer Congress, San Francisco 1989, North-Holland, pp. 613–618. Available

at http://theory.stanford.edu/~rvg/abstracts.html#11. Full version in Jounal of the

ACM 43(3), 1996, pp. 555–600.

[206] R.J. van Glabbeek & W.P. Weijland (1989): Refinement in Branching Time Semantics. Report CS-

R8922, CWI, Amsterdam. Available at http://theory.stanford.edu/~rvg/abstracts.

html#12. Also appeared in: Proceedings AMAST Conference, May 1989, Iowa, USA, pp. 197– 201.

[207] R.J. van Glabbeek & W.P. Weijland (1989): Refinement in Branching Time Semantics. In: Proceedings AMAST Conference, May 1989, Iowa, USA, pp. 197–201. Available at http://

theory.stanford.edu/~rvg/abstracts.html#12.

[208] R.J. van Glabbeek & W.P. Weijland (1990): Branching Time and Abstraction in Bisimulation Semantics. Technical Report TUM-I9052, SFB-Bericht Nr. 342/29/90 A, Institut für Informatik, Technische Universität München, Germany. Available at http://theory.stanford. edu/~rvg/abstracts.html#23. Original version of [209]. Extended abstract in G.X. Ritter, editor: Information Processing 89, Proceedings of the IFIP 11th World Computer Congress, San Francisco, USA 1989, Elsevier Science Publishers B.V. (North-Holland), 1989, pp. 613–618.

[209] R.J. van Glabbeek & W.P. Weijland (1996): Branching Time and Abstraction in Bisimulation Semantics. Journal of the ACM 43(3), pp. 555–600, doi:10.1145/233551.233556. Available in

part at http://Theory.Stanford.EDU/~rvg/abstraction/.

[210] P. Höfner, R.J. van Glabbeek, W.L. Tan, M. Portmann, A.K. McIver & A. Fehnker (2012): A Rigorous Analysis of AODV and its Variants. In A.Y. Zomaya, B. Landfeldt & R. Prakash, editors: Proceedings 15th ACM International Conference on Modeling, Analysis and Simulation of Wireless and Mobile Systems, MSWiM ’12, Paphos, Cyprus, October 2012, ACM, pp. 203–

212, doi:10.1145/2387238.2387274. Available at http://theory.stanford.edu/~rvg/
abstracts.html#97.

[211] P. Höfner, R.J. van Glabbeek & I.J. Hayes (2012): Preface—Morgan: a suitable case for treatment. Formal Aspects of Computing 24(4–6), pp. 417–422, doi:10.1007/s00165-012-0257-0.

Available at http://theory.stanford.edu/~rvg/abstracts.html#96. Festschrift, Cele-

brating the 60th Birthday of Carroll Morgan.

[212] D.J.D. Hughes & R.J. van Glabbeek (2003): Proof Nets for Unit-free Multiplicative-Additive Linear Logic (extended abstract). In: Proceedings 18^{th} Annual IEEE Symposium on Logic in Computer Science, LICS’03, Ottawa, Canada, June 2003, IEEE Computer Society Press, pp. 1–

10, doi:10.1109/LICS.2003.1210039. Available at http://theory.stanford.edu/~rvg/
abstracts.html#50.

[213] D.J.D. Hughes & R.J. van Glabbeek (2005): Proof Nets for Unit-free Multiplicative-Additive Linear Logic. ACM Transactions on Computational Logic 6(4), pp. 784–842, doi:10.1145/

1094622.1094629. Available at http://theory.stanford.edu/~rvg/abstracts.html#
57.

[214] C. Lippert, S. Mennicke, R.J. van Glabbeek & U. Goltz (2018): A Case Study on Evaluating Encodings Between Process Calculi. Technical Report, Data61, CSIRO. Available at http://

theory.stanford.edu/~rvg/abstracts.html#129.

[215] E.-R. Olderog, U. Goltz & R.J. van Glabbeek (1988): Combining Compositionality and Concurrency, Summary of a GMD-Workshop, Königswinter, March 1988. Arbeitspapiere der GMD 320, Gesellschaft für Mathematik und Datenverarbeitung, Sankt Augustin.

[216] C. Palamidessi, J. Parrow & R.J. van Glabbeek (2000): Preface to the special issue of Information and Computation dedicated to EXPRESS’97. Information and Computation 156, p. 1, doi:10.

1006/inco.1999.2815.

[217] K. Peters & R.J. van Glabbeek (2015): Analysing and Comparing Encodability Criteria. In S. Crafa & D.E. Gebler, editors: Proceedings of the Combined 22th International Workshop on Expressiveness in Concurrency and 12th Workshop on Structural Operational Semantics, Madrid, Spain, 31st August 2015, Electronic Proceedings in Theoretical Computer Science 190, Open Publishing Association, pp. 46–60, doi:10.4204/EPTCS.190.4.

[218] K. Peters & R.J. van Glabbeek (2015): Analysing and Comparing Encodability Criteria for Process Calculi. Archive of Formal Proofs 2015. Available at http://afp.sourceforge.net/

entries/Encodability_Process_Calculi.shtml.

[219] X. Qin, L. O’Connor, R.J. van Glabbeek, P. Höfner, O. Kammar & M. Steuwer (2024): Shoggoth: A Formal Foundation for Strategic Rewriting. Proceedings of the ACM on Programming Languages 8(POPL), pp. 61–89, doi:10.1145/3633211.

[220] P.H. Rodenburg & R.J. van Glabbeek (1988): An interpolation theorem in equational logic. Report CS-R8838, CWI, Amsterdam. Available at http://theory.stanford.edu/~rvg/

abstracts.html#9.

[221] D.G. Stork & R.J. van Glabbeek (2002): Token-controlled place refinement in hierarchical Petri nets with application to active document workflow. In J. Esparza & C. Lakos, editors: Proceedings 23^{rd} International Conference on Application and Theory of Petri Nets, ICATPN’02, Adelaide, Australia, June 2002, LNCS 2360, Springer, pp. 394–413, doi:10.1007/3-540-48068-4_23.

Available at http://theory.stanford.edu/~rvg/abstracts.html#47.